# Linear Model of Aggregated Homogeneous Energy Storage Elements with Realizable Dispatch Guarantees

### v4

@@ III-D Realizable Composite Power Trajectories
Note that for a fixed composite battery, if $N$ increases, the cutting plane in (13c) approaches the cutting plane in (4b). Furthermore, for $M\gg 1$, $\varepsilon$ in (10) tends to zero. That is, with faster control and more granular elements, (13) approaches the relaxed battery model defined in (4); however, with guaranteed realizability.
The energy buffer $\varepsilon$ in (13d) restricts the set of feasible energy trajectories relative to (3) and (4a). The feasible energy trajectories in the robust linear model presented in [RobustNawaf] are also a subset of the feasible energy trajectories in (3) and (4a). However, the proposed model in (13) has a constant energy buffer $\varepsilon$ that is independent of the length of time horizon considered. This is in contrast to the robust model which has an energy buffer that increases with the length of the time horizon which, in the worst case, significantly reduces the feasible set of energy trajectories and can yield sub-optimal solutions. This is illustrated in Fig. 4 where the (worst case) feasible set of energy trajectories for the robust model and the RCB model for $\Delta t=1\text{hr}$ and $M=1$ ($\delta t=1\text{hr}$) and $M=10$ ($\delta t=6\text{min}$) are compared for the battery parameters of a Tesla Powerwall shown in Table I.Clearly, smaller control time steps result in a smaller $\varepsilon$ and thus a larger feasible set of energy trajectories for the proposed model, while the robust model remains unchanged.
@@ IV Simulation Results
We apply the realizable composite battery (RCB) model in (13) to two relevant use cases and compare its performance against the tighter relaxed model given in (4) and the robust model in [RobustNawaf] used under balanced operation. Also, a Mixed-Integer Linear Program (MILP) formulation with unequal power-sharing for $N$ elements in (1) is used for comparison. The battery element parameters given in Table I were used. The two use-cases are a) power reference tracking and b) revenue maximization.
@@ IV-A Power Reference Tracking
A timestep of $\Delta t=5$ min, along with $N=100$ was used. The RCB model was used three times with $M=1$, $M=5$, and $M=10$ respectively. The results from using each model are shown in Fig. 5 and the solve times along with the mean squared tracking error are shown in Table II.
@@
TABLE II: Power Reference Tracking Performance Metrics
@@
Time (ms) MSE (kW2) MSE (kW2)
MILP 50,140 0 0
Relaxed 30 0 1841
Robust 34 2.12 2.12
RCB (M = 1) 30 0.72 0.72
RCB (M = 5) 30 0.013 0.013
RCB (M = 10) 30 0 0
The MILP model yielded the optimal solution and was able to track the reference signal with no tracking error. However, the MILP also came with a significant computational burden compared to the rest of the formulations. The other formulations are linear and, as a result, have significantly faster solve times. The relaxed model predicts zero tracking error; however, the solution was unrealizable and caused the battery to saturate around $t=6$ hrs leading to a large realized tracking error. The robust model yields a realizable yet sub-optimal solution. The RCB model shows improved tracking error with increasing values of $M$ but even with $M=1$ the tracking error surpasses the robust model.
@@ IV-B Revenue Maximization
where $C[k]$ is the price at time $k$. The price signal was obtained from the California Independent System Operator (CAISO) day-ahead market prices for 24 hours, plotted in Fig 6.
@@
TABLE III: Revenue Maximization Performance metrics
Model Solver Predicted Actual
Time (ms) Revenue (\$) Revenue (\$)
MILP 1250 2088.97 2088.97
Relaxed 1 2092.80 1940.35
Robust 1 1866.42 1866.42
RCB (M = 1) 1 1411.53 1411.53
RCB (M = 5) 1 1955.76 1955.76
RCB (M = 10) 1 2022.87 2022.87
RCB (M = 900) 1 2089.23 2089.23
Similar to the previous use-case, the MILP model produced the highest revenue but at a significantly larger computational cost. The relaxed model predicts the highest revenue but underestimates the SOC, leading to unpredicted battery saturation and a realized performance that is significantly worse. The robust model yields a realizable yet suboptimal solution. The RCB model for small values of $M$ results in a sub-optimal solution which can be attributed to the larger timestep used in this use-case leading to a larger value or $\varepsilon$ and, in turn, a more restricted feasible set. However, for f $M\geq 5$ ($\delta t\leq 3$ min), the RCB model surpasses all other linear models and approaches the predicted performance for larger values of $M$. It was also observed that the RCB surpasses the robust model for $M\geq 3$ ($\delta t\leq 5$ min).
@@ V Conclusion
A new linear dispatch model is presented and analyzed for a composite battery that permits unequal power sharing among its $N$ elements to realize simultaneous charging and discharging. A priority-based disaggregation strategy is described and used to characterize sufficient conditions on the composite battery dispatch under which element constraints are guaranteed to be satisfied. Previous battery optimization implicitly assumed equal power sharing between all elements. By enabling unequal power sharing within a linear formulation, we have expanded the class of battery models that yield realizable power trajectories. Practical implementation of such dispatch strategies could also justify the use of linear models for composite storage systems or aggregate storage representations in energy system optimization models.

### v6

@@ III-D Realizable Composite Power Trajectories
Note that for a fixed composite battery, if $N$ increases, the cutting plane in (13c) approaches the cutting plane in (4b). Furthermore, for $M\gg 1$, $\varepsilon$ in (10) tends to zero. That is, with faster control and more granular elements, (13) approaches the relaxed model in (4); however, with guaranteed realizability.
The energy buffer $\varepsilon$ in (13d) restricts the set of feasible energy trajectories relative to (3) and (4a). The feasible energy trajectories in the robust linear model presented in [RobustNawaf] are also a subset of the feasible energy trajectories in (3) and (4a). However, the proposed model in (13) has a constant energy buffer $\varepsilon$ that is independent of the length of time horizon considered. This is in contrast to the robust model which has an energy buffer that increases with the length of the time horizon which, in the worst case, significantly reduces the feasible set of energy trajectories and can yield sub-optimal solutions. This is illustrated in Fig. 3 where the (worst case) feasible set of energy trajectories for the robust model and the RCB model for $\Delta t=1\text{hr}$ and $M=1$ ($\delta t=1\text{hr}$) and $M=5$ ($\delta t=12\text{min}$) are compared for the battery parameters of a Tesla Powerwall shown in Table I. Clearly, smaller control time steps result in a smaller $\varepsilon$ and thus a larger feasible set of energy trajectories for the proposed model, while the robust model remains unchanged.
@@ IV Simulation Results
We apply the realizable composite battery (RCB) model in (13) to two relevant use cases and compare its performance against the tighter relaxed model given in (4) and the robust model in [RobustNawaf] used under balanced operation. Also, a Mixed-Integer Linear Program (MILP) formulation with both equal and unequal power-sharing for $N$ elements in (1) used for comparison. The battery element parameters given in Table I were used. The two use-cases are a) power reference tracking and b) revenue maximization. All cases are modeled using JuMP in Julia and were solved using Gurobi v12.0.0. All cases were run on a Macbook Pro with an M1 Pro processor.
@@ IV-A Power Reference Tracking
A timestep of $\Delta t=3$ min, along with $N=100$ was used. The RCB model was used three times with $M=1$, $M=5$, and $M=10$ respectively. The results from using each model are shown in Fig. 4 and the solve times along with the mean squared error (MSE) are shown in Table II. Note that the relaxed model resulted in a dispatch with simultaneous charging and discharging, which cannot be realized under equal power sharing and results in the discrepancy between predicted MSE and actual MSE obtained while realizing the optimal dispatch. The net power trajectory was implemented which caused the BESS to saturate causing the discrepancy between the predicted and realized MSE values.
@@
TABLE II: Power Reference Tracking Performance Metrics ($N=100$)
@@
Time MSE (kW2) MSE (kW2)
MILP (unequal) 492.60 s 0 0
MILP (equal) 34.37 s 200.0 200.0
Relaxed 45 ms 0 1889
Robust 54 ms 212 212
RCB (M = 1) 50 ms 23.93 23.93
RCB (M = 5) 50 ms 0.231 0.231
RCB (M = 10) 50 ms 0 0
The MILP model under unequal power sharing was able to track the reference signal with no tracking error. The MILP under equal power sharing couldn’t track the reference signal as well, despite both solving to 0% optimality gap, indicating that there is a benefit of simultaneously charging and discharging in this use case. The MILP, under both equal and unequal power sharing, came with a significant computational burden compared to the rest of the formulations, attributed to the extra binary variables added to enforce the complementarity constraint. The other formulations are linear and, as a result, have significantly faster solve times. The relaxed model predicts zero tracking error; however, the solution was unrealizable and caused the battery to saturate around $t=6$ hrs, leading to a large realized tracking error. The robust model yields a realizable yet sub-optimal solution. The RCB model shows improved tracking error with increasing values of $M$, but even surpassing the robust model with $M=1$.
@@ IV-B Revenue Maximization
where $C[k]$ is the price at time $k$. The price signal was obtained from the California Independent System Operator (CAISO) day-ahead market prices, plotted in Fig 5.
@@
TABLE III: Revenue Maximization Performance metrics ($N=100$)
Model Solver Predicted Actual Optimality
Time Revenue (\$) Revenue (\$) Gap (%)
MILP (unequal) $90$ mins 2088.99 2088.99 0.02
MILP (equal) 23 mins 2088.97 2088.97 0
Relaxed 1.5 ms 2092.80 1940.35 0
Robust 2.2 ms 1866.42 1866.42 0
RCB (M = 1) 1.7 ms 1642.24 1642.24 0
RCB (M = 5) 1.9 ms 2000.50 2000.50 0
RCB (M = 10) 1.7 ms 2045.24 2045.24 0
RCB (M = 900) 1.6 ms 2089.48 2089.48 0
The MILP models produced high revenues that could also be realized, but at substantial computational cost, with unequal power sharing taking significantly longer to solve and was not able to solve to 0% optimality gap. The relaxed model predicts the highest revenue but underestimates the SOE, leading to unpredicted battery saturation and a realized performance that is significantly worse. The robust model yields a realizable yet suboptimal solution. The RCB model for small values of $M$ results in a sub-optimal solution which can be attributed to the larger timestep used in this use-case leading to a larger value or $\varepsilon$ and, in turn, a more restricted feasible set. However, for $M\geq 5$ ($\delta t\leq 3$ min), the RCB model surpasses all other linear models and for larger values of $M$ the predicted revenue steadily increases, maintaining realizability.
We note that for $M=900$ the revenue even surpasses the element-wise MILP solution, because individual elements are able to switch from charging to discharging and back at a faster controller timestep when implemented by the PSC. A MILP solution with the control time step $\delta t$ would again achieve or exceed this result, but at much larger computational cost. The RCB model captures what is feasible at the fast control time step $\delta t$ without added complexity.
@@ V Conclusion
A new linear dispatch model is presented and analyzed for a composite battery that permits unequal power sharing among its $N$ elements. A priority-based disaggregation strategy is described and used to characterize sufficient conditions on the composite battery dispatch under which element constraints are guaranteed to be satisfied. Previous battery optimization implicitly assumed equal power sharing between all elements. By enabling unequal power sharing within a linear formulation, we have expanded the class of battery models that yield realizable power trajectories. Implementing the proposed controller is of interest and mainly requires communications to/from battery elements faster than $\delta t$ (similar to VPPs in the field today). Furthermore, embedding the presented linear models into energy system planning problems would inform the impact of battery systems more broadly. Future work involves exploring other disaggregation strategies and non-linear battery models.